OpenStudy (anonymous):

If the circumference of a circle is 12, and the angle measure of an arc is 60o, which is the length of the arc? 2 3 6 12

OpenStudy (igreen):

Find the radius using the formula to find the circumference of a circle. C = 2pi r 12 = 2(3.14)r Solve for 'r', can you do that?

OpenStudy (igreen):

Oh wait..the equation to find the arc length is: \(\sf 2 \pi r(\dfrac{\theta}{360})\)

OpenStudy (igreen):

2pi r is the same as the circumference..which in this case is 12. So we have: \(\sf 12(\dfrac{\theta}{360})\) And the angle measurement here is 60, so plug it in and simplify, can you do that?

OpenStudy (anonymous):

Thanks man

OpenStudy (igreen):

No problem.